The average marks of boys in class is 52 and that of girls is 42. The average marks of boys and girls combined is 50. The percentage of boys in the class is ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Let the number of boys and girls be x and y.
52x+42y=50(x+y)
x=4y
Total number of students in the class=x+y=5y
Required % of boys=(4y/5y)100%=80%
Thanks and Regards
Shaik Aasif
askIITians faculty